OpenStudy (anonymous):

What is the solution of log(2x – 6) 256 = 4? I just want to know how to do this.

OpenStudy (jhannybean):

is 256 a power or multiplier?

OpenStudy (anonymous):

multiplier

OpenStudy (jhannybean):

So your problem is: \(256\log(2x-6)=4\) ?

OpenStudy (anonymous):

Oh wait xD... my problem is like...

OpenStudy (anonymous):

\[\log_{2x-6} 256 = 4\]

OpenStudy (jhannybean):

okay. remember the log rule: \(\large \log_x b=a \implies x^a=b\)

OpenStudy (anonymous):

I have no idea what that means xD

OpenStudy (jhannybean):

\[\log_{(2x-6)}256 = 4 \implies (2x-6)^4 = 256\] Do you see how the formula works?

OpenStudy (anonymous):

Yes, but how do I change that to the answer? (The answer is x = 5, I took the test, I just wanna know how to do it)

OpenStudy (jhannybean):

Alright

OpenStudy (jhannybean):

to get rid of the 4rth power, you take the 4rth root from both sides o the equation.

OpenStudy (jhannybean):

\[(2x-6)^4=256\]\[\sqrt[4]{(2x-6)^4} = \sqrt[4]{256}\]\[2x-6=4\]\[2x=10\]\[x=5\]

OpenStudy (anonymous):

Thank you, thats all i wanted haha, I already gave u a medal